# Let it Snow! Animating 3D Gaussian Scenes with Dynamic Weather Effects via Physics-Guided Score Distillation

### v2

@@ B.1 Implementation Details
We then add particles based on their effect-specific motion simulation parameters (detailed in Sec. B.3). We run the simulation for 500 steps. We set the threshold for particle active state updates at 0.001. Our grid resolution is set to $64^{3}$. After each simulation step we transform the positions back to the Gaussian world and save them for rendering. We also save the collided Gaussians position at the time of collision.
@@
Runtime. Our framework runs on a single NVIDIA A100 40GB GPU. Physics simulation takes approximately 2 minutes and physics-guided optimization takes approximately 1 hour. Once optimized, dynamic effect rendering achieves 10 FPS.
B.2 Collision Handling Details
@@
B.3 Motion Simulation Parameters
@@
B.4 Appearance Initialization
We detail our appearance parameter LLM-initialization. We employ a large language model (Claude Sonnet 4, Anthropic, 2025) to generate baseline parameters $A^{active}_{init}$ and $A^{collided}_{init}$ for each effect. We prompt the LLM with detailed descriptions to obtain appropriate scale, opacity, and color parameters.
@@ B.4 Appearance Initialization
B.5 Comparisons and Ablation Details
@@
w/o Collision Handling (w/o CH). We disable the mesh-based collision handling techniques described in Sec. B.2. Dynamic Gaussians follow their simulated trajectories without surface projection or accumulation adjustments, resulting in Gaussians floating above surfaces rather than realistically interacting with scene geometry.
w/o Appearance Optimization (w/o App). We skip the Video-SDS appearance optimization stage entirely and directly use the LLM-initialized parameters described in Sec. B.4. This variant uses only the baseline appearance parameters $\mathcal{A}^{active}_{init}$ and $\mathcal{A}^{collided}_{init}$ without any refinement through diffusion guidance.
@@
Appendix C Additional Visualizations and Results

### v4

@@ 8.1 Implementation Details
We then add particles based on their effect-specific motion simulation parameters (detailed in Sec. 8.3). We run the simulation for 500 steps. We set the threshold for particle active state updates at 0.001. Our grid resolution is set to $64^{3}$. After each simulation step we transform the positions back to the Gaussian world and save them for rendering. We also save the collided Gaussians position at the time of collision.
@@
Runtime. Our framework runs on a single NVIDIA A100 40GB GPU. A runtime comparison against baselines is provided in Tab. 11.
Table 11: Runtime Comparison. Per scene-effect runtime on a single NVIDIA A100 40GB GPU.
Method Runtime
ClimateNeRF [18] $\sim$}180 min
GaussCtrl [36] $\sim$}30 min
Instruct-4DGS [16] $\sim$}45 min
Ours $\sim$}60 min
8.2 Collision Handling Details
Following Sec. 3.2, we describe our mesh-based collision handling.
@@ 8.2 Collision Handling Details
8.3 Motion Simulation Parameters
Following Sec. 3.2, we detail our motion simulation parameters.
@@ 8.3 Motion Simulation Parameters
8.4 Appearance Initialization
Following Sec. 3.3, we detail our appearance parameter LLM-initialization. We employ a large language model (Claude Sonnet 4, Anthropic, 2025) to generate baseline parameters $A^{active}_{init}$ and $A^{collided}_{init}$ for each effect. We prompt the LLM with detailed descriptions to obtain appropriate scale, opacity, and color parameters.
@@ 8.4 Appearance Initialization
8.5 Comparisons and Ablation Details
@@
w/o Collision Handling (w/o CH). We disable the mesh-based collision handling techniques described in Sec. 8.2. Dynamic Gaussians follow their simulated trajectories without surface projection or accumulation adjustments, resulting in Gaussians floating above surfaces rather than realistically interacting with scene geometry.
w/o Appearance Optimization (w/o App). We skip the Video-SDS appearance optimization stage entirely and directly use the LLM-initialized parameters described in Sec. 8.4. This variant uses only the baseline appearance parameters $\mathcal{A}^{active}_{init}$ and $\mathcal{A}^{collided}_{init}$ without any refinement through diffusion guidance.
@@
9 Additional Visualizations and Results